OpenStudy (anonymous):

determine if the statement is true or false the slope-intercept form of -4y-7=5 is y= -5/4x-7/4

OpenStudy (anonymous):

my dear once again there is a missing \(x\)

OpenStudy (anonymous):

could it be \(-4y-7x=5\)?

OpenStudy (anonymous):

or maybe \(-4y-7=5x\) ?

OpenStudy (anonymous):

ops sorry wrong equation

OpenStudy (anonymous):

lol i will wait while you change it

OpenStudy (anonymous):

\[-4y-7=5x\]

OpenStudy (anonymous):

sorry xD

OpenStudy (anonymous):

and you want the slope, is that right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok i made a typo there \[-4y-7=5x\] solve in two steps add \(7\) to both sides to get \(-4y=5x+7\) divide both sides by \(-4\) to get \(y=-\frac{5}{4}x-\frac{7}{4}\)

OpenStudy (anonymous):

ok skip that , it is just True of False right? looks True to me

OpenStudy (anonymous):

yep its true !

OpenStudy (anonymous):

you got this?

OpenStudy (anonymous):

yes